A sample size 25 is picked up at random from a population which is normally distributed with a mean of 100 and variance of 36. Calculate:-

(a) Pr{ ͞X≤99}

(b) Pr{98≤ ͞X≤100}


1
Expert's answer
2021-05-07T10:40:27-0400

"n = 25 \\

\mu=100 \\

\sigma^ = 36 \\
\sigma = 6"

(a)

"z = \frac{\bar{X}- \mu}{\sigma / \sqrt{n}} \\

P(\bar{X} ≤ 99) = P(Z≤ \frac{99 -100}{6/ \sqrt{25}} )\\

= P(Z≤ -0.8333) \\

= 0.2024"

(b)

"P(98≤\bar{X} ≤ 100) = P(\bar{X}≤100) -P(\bar{X}≤98) \\

= P(Z≤ \frac{100-100}{6/ \sqrt{25}}) -P(Z≤ \frac{98-100}{6/ \sqrt{25}} )\\

= P(Z≤ 0) -P(Z≤ -1.666) \\

= 0.5 -0.0478 \\

= 0.4522"